OpenStudy (anonymous):

MEDAL! HELP ME NOW OR I WILL SURELY DIE!!!!! Helen has been working on a science project about plant growth. Plant A started at a height of 6 cm and exhibited steady growth during the first four weeks of the project. After those four weeks, Plant A was 26.8 cm in height. In a linear model of Plant A's growth, x represents the number of weeks after being planted, and y represents the plant's height [in cm]. Relative to this model, which of the following statements applies? A. An additional week of growth is associated with an additional 8.2 cm of plant height. B. An additional 8.2 weeks of growth is associated with an additional centimeter of plant height. C. An additional 5.2 weeks of growth is associated with an additional centimeter of plant height. D. An additional week of growth is associated with an additional 5.2 cm of plant height.

OpenStudy (anonymous):

@Data_LG2 ?

OpenStudy (anonymous):

let's analyse the problem first we need this: y=mx+b original height will be 6 (y-intercept-b) x will be the time ( m (slope) is the growth rate of the plant so far you have: y= mx + 6 you are given: after 4 weeks y=26.8 x = 4 now let's for the growth rate or m using the given values 26.8 = m (4) + 6 can you simplify this formula for me? i'll brb

OpenStudy (anonymous):

Okay, thanks! I'll simplify it.

OpenStudy (anonymous):

m=5.2. So now I have to figure out if it's C or D.....right?

OpenStudy (anonymous):

Nvm, the question expired, Thanks for helping though.

OpenStudy (anonymous):

i'm back, i'm sorry what do you mean by the question is already expired? O.o anyway, if in case you're still interested to know how to get the answer.. y=5.2 x + 6 (so you have this eqn) let's analyse each choices now. Choice A: An additional week of growth is associated with an additional 8.2 cm of plant height. This means when x=5, y should be 26.8 + 8.2 = 35 let's solve y: y =5.2 (5) + 6 = 32 cm, therefore this choice is not right Choice B: An additional 8.2 weeks of growth is associated with an additional centimeter of plant height. This means when x=8.2+4= 12.2 , y should be 26.8 + 1 = 27.8 let's solve y: y =5.2 (8.2) + 6 = 48.64 cm, therefore this choice is not right Choice C: An additional 5.2 weeks of growth is associated with an additional centimeter of plant height. This means when x=5.2+4= 9.2 , y should be 26.8 + 1 = 27.8 let's solve y: y =5.2 (9.2) + 6 = 53.84 cm, therefore this choice is not right D. An additional week of growth is associated with an additional 5.2 cm of plant height. This means when x=4+1= 5, y should be 26.8 + 5.2 = 32 let's solve y: y =5.2 (5) + 6 = 32 cm, therefore this choice is the right choice (:

OpenStudy (anonymous):

Thanks a lot! When I said the question expired I meaant I was on Study Island and becasue I was gone from it to long, it just quit the entire session. It's not your fault tho. I still needed to know how to get the answer for future refrence so thanks a bunch!

OpenStudy (anonymous):

oh sorry to hear that :/ goodluck with your studies (:

OpenStudy (anonymous):

Thanks! And an extra thanks 4 the medal! :)